OpenStudy (gabylovesyou):

The graph below shows the revenue earned by a company in its first four quarters of operation. A line of best fit has been drawn on the graph.

OpenStudy (gabylovesyou):

Based on the linear model, how much revenue, in thousands of dollars, should the company earn in its 8th quarter of operation? Type your answer in the box below.

OpenStudy (gabylovesyou):

@phi @ganeshie8 @ParthKohli

OpenStudy (phi):

you should estimate the slope of this line.

OpenStudy (gabylovesyou):

well... umm i really have no clue :S

OpenStudy (phi):

it looks like about 10 next, pick a point on the line, like (3,30) find the equation of the line: y - 30 = 10( x-3) or y-30 = 10x -30 y = 10x check, to make sure it makes sense: with x=1, y =10 which is a decent match

OpenStudy (phi):

to find the slope, start at x=3, y=30 then count up 10 and over 1 the change is y is +10 and the change in x is +1 and the slope is 10/1 = 10

OpenStudy (phi):

how much does it earn in its 8th quarter of operation? use y=10x with x=8 to find y=80

OpenStudy (gabylovesyou):

yup :p

OpenStudy (gabylovesyou):

so its gonna be what (8,80) ???

OpenStudy (gabylovesyou):

or 80 for the 8th thing ?

OpenStudy (phi):

I think they just want the y value (just the earnings)

OpenStudy (gabylovesyou):

ohh ok ! thank you!!!!